OpenStudy (anonymous):

Isabel is 11 years old. This is 3 years more than twice her brother's age. How old is Isabel's brother? Let b = Isabel's brother's age. A. 3 - 2b = 11; b = 16 years old B. 2(11 - 3) = b; b = 16 years old C. 2b + 3 = 11; b = 4 years old D. 2b - 3 = 11; b = 7 years old

OpenStudy (anonymous):

Isabel is 11 yrs old. Let Isabel's age =x then we are told this is 3 more years than twice her brothers age. If her brother age is b, then that equates to 2b -3 i.e x (her age)= 2b -3

OpenStudy (anonymous):

since she is 11, 2b-3 =11 and 2b = 11 +3 = 14 b = 14/2 which equals 7 therefore her brother is 7

OpenStudy (anonymous):

answer is D